# The Unbounded Denominators Conjecture

### v2

@@ 1. Introduction
We refer the reader to Atkin and Swinnerton-Dyer [ASD71] for the roots of the unbounded denominators conjecture, and to Birch’s article [Bir94] as well as to Long’s survey [Lon08, § 5] for an introduction to this problem and its history. For the vector-valued generalization, see §7.3 and its references below. The cases of relevance to the partition and correlation functions of rational conformal field theories (of which the tip of the iceberg is the example (1.0.3) discussed below) were resolved in a string of works [DR18, DLN15, SZ12, NS10, Xu06, Ban03, Zhu96, AM88], by the modular tensor categories method. Some further sporadic cases of the unbounded denominators conjecture have been settled by mostly ad hoc means [FF20, FM16b, LL12, KL08, KL09].
@@ 1.1.3. A finite local monodromy leads to an overconvergence
But if instead of the tautological simultaneous uniformization we take $\varphi:D(0,1)\to\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ to be the universal covering map (pointed at $\varphi(0)=0$), then either by a direct computation with monodromy, or by Cauchy’s analyticity theorem on the solutions of linear ODEs with analytic coefficients and no singularities in a disc, we have both function germs $x:=\varphi(z)$ and $f(x):=f(\varphi(z))$ holomorphic, hence convergent, on the full unit disc $D(0,1)$. In contrast, now $F(\varphi(z))$ converges only up to the “first” nonzero fiber point in $\varphi^{-1}(0)\smallsetminus\{0\}$, giving a certain radius rather smaller than $1$. We must have the strict lower bound $|\varphi^{\prime}(0)|>1$, because the preceding unit-radius holomorphic map $\sqrt[N]{\lambda(z^{N})/16}:D(0,1)\to\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ has to factorize properly through the universal covering map. Indeed in Theorem 5.1.4, using an explicit description by hypergeometric functions of the multivalued inverse of the universal covering map of $\mathbf{C}\smallsetminus\mu_{N}$ based on Poincaré’s ODE approach [Hem88] to the uniformization of Riemann surfaces, we find an exact formula for this uniformization radius in terms of the Euler Gamma function. [^2] André pointed out to us that this explicit formula has previously been obtained by Kraus and Roth, see [KR16, Remark 5.1]. See also [Gol69, §III.1]. Hence the algebraicity of $f(x)$ gets witnessed by André’s criterion; and the formal new result that we get already at this opening stage (see Theorem 7.2.1) is that any integral formal power series solution $f(x)\in\mathbf{Z}\llbracket x\rrbracket$ to a linear ODE $L(f)=0$ without singularities on $\mathbf{P}^{1}\smallsetminus\{0,1/16,\infty\}$ is in fact algebraic as soon as the linear differential operator $L$ has a finite local monodromy $\mathbf{Z}/N$ around the singular point $x=0$. More than this: the quantitative Corollary 2.0.5 proves that the totality of such $f(x)\in\mathbf{Z}\llbracket x\rrbracket$ at a given $N$ span a finite-dimensional $\mathbf{Q}(x)$-vector space, and gives an upper bound on its dimension as a function of the Wohlfahrt level parameter $N$. Now since a (noncongruence) counterexample $f(\tau)\in\mathbf{Z}\llbracket q\rrbracket$ to Theorem 1.0.1 would not exist on its own but spawn a whole sequence $f(\tau/p)\in\mathbf{Z}\llbracket q^{1/p}\rrbracket$ of $\mathbf{Q}(x)$-linearly independent counterexamples at growing Wohlfahrt level $N\mapsto Np$, our idea is to measure up the supply of these putative (fictional) counterexamples alongside the congruence supply at a gradually increasing level until together they break the quantitative bound (2.0.3) supplied by our arithmetic holonomy Theorem 2.0.1.
@@ 1.1.4. The dimension bound can be leveraged with growing level NN
We have the congruence supply of dimension $\Gamma(2):\Gamma(2N)]\gg N^{3}$, and then as a glance at our shape (2.0.7) of holonomy rank bound readily reveals, it seems a fortuitous piece of luck that the conformal size (Riemann uniformization radius at 0) of our relevant Riemann surface $\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ turns out to have the matching asymptotic form $1+\zeta(3)/(2N^{3})+O(N^{-5})$. We “only” have to prove that the numerator (growth) term in the holonomy rank bound (2.0.7) inflates at a slower rate than our extrapolating putative counterexamples $f(\tau)\mapsto f(\tau/p)$!
@@ 2. The arithmetic holonomicity theorem
We invite the reader on a first pass to proceed directly to §3 after §2.3.
@@ 2.1. The auxiliary construction
2.2. Extrapolation and proof of Theorem 2.0.8
@@
$\sup_{\mathbf{T}^{d}}\log{|H|}\leq dD\cdot\max\big\{\sup_{\mathbf{T}}\log{|h|},\,\sup_{\mathbf{T}}\log{|h\cdot\varphi^{*}p|}\big\}+\kappa C\alpha+o(\alpha)$
$\leq\Big(\kappa C+\frac{d}{(d!)^{1/d}}\Big(1+\frac{1}{\kappa}\Big)^{\frac{1}{d}}\Big)\cdot\frac{\alpha}{m}+o(\alpha).$
@@
$\alpha\log{|\varphi^{\prime}(0)|}\leq\Big(\kappa C+\frac{d}{(d!)^{1/d}}\Big(1+\frac{1}{\kappa}\Big)^{\frac{1}{d}}\Big)\cdot\frac{\alpha}{m}\max\big\{\sup_{\mathbf{T}}\log{|h|},\,\sup_{\mathbf{T}}\log{|h\cdot\varphi^{*}p|}\big\}+o(\alpha)$
@@
$m\leq\inf_{\kappa\in(0,1),\,d\in\mathbf{N}}\Big\{\Big(\kappa C+\frac{d}{(d!)^{1/d}}\Big(1+\frac{1}{\kappa}\Big)^{\frac{1}{d}}\Big)\Big\}\cdot\frac{\max\big\{\sup_{\mathbf{T}}\log{|h|},\,\sup_{\mathbf{T}}\log{|h\cdot\varphi^{*}p|}\big\}}{\log{|\varphi^{\prime}(0)|}}.$
Theorem 2.0.8 now follows by firstly letting $d\to\infty$ and then $\kappa\to 0$, and observing that in that limit
$\kappa C+\frac{d}{(d!)^{1/d}}\Big(1+\frac{1}{\kappa}\Big)^{\frac{1}{d}}\to e,$
@@
At this point Theorem 2.0.1 comes as the immediate combination of Theorem 2.0.8 and the following classical lemma of Nevanlinna.
@@
Every holomorphic function $g:\overline{D(0,1)}\to\mathbf{C}$ admits a quotient representation
@@
$h(0)=1\quad\textrm{and}\quad\max\big\{\sup_{\mathbf{T}}\log{|h|},\,\sup_{\mathbf{T}}\log{|hg|}\big\}\leq\int_{\mathbf{T}}\log^{+}{|g|}\,\mu_{\mathrm{Haar}}.$
@@ 2.5.4. A potential-theoretic generalization
3. Some reductions
The following encapsulates our argument:
@@ 3. Some reductions
We may also prove this proposition by using Theorem 2.0.1 directly. Using the notation in the proof, let $Y^{\prime}$ denote the modular curve $Y$ with all the cusps above $0\in Y(2)$ filled in. The fiber product $Y^{\prime}\times_{Y(2)_{\overline{\mathbf{Q}}}}U$ with its natural map to $U$ is a covering map (one can check this claim locally; the assumption on cusp widths is used to prove that $0\in U$ is not ramified). Therefore, the universal covering map $D(0,1)\rightarrow U$ factors through $Y^{\prime}\times_{Y(2)_{\overline{\mathbf{Q}}}}U$ and thus we obtain a map $D(0,1)\rightarrow Y^{\prime}\times_{Y(2)_{\overline{\mathbf{Q}}}}U\rightarrow Y^{\prime}$ (the second map is the natural map) such that its composition with $Y^{\prime}\rightarrow Y(2)_{\overline{\mathbf{Q}}}\cup\{0\}$ is the map $D(0,1)\rightarrow U\rightarrow Y(2)_{\overline{\mathbf{Q}}}\cup\{0\}$. Thus $f\circ\varphi$ is also given by the natural pullback of $f$ from $Y^{\prime}$ to $D(0,1)$ and thus it is holomorphic over $D(0,1)$ as far as $f$ is holomorphic at all cusps in $Y^{\prime}$, which can be achieved by multiplying $f$ with a suitable power of $\lambda$. Thus we verify the analyticity property in Theorem 2.0.1. The rest of the proof is the same as the proof above.
We prove both of the required statements in Theorems 5.1.4 and Theorem 6.0.1 respectively. We then put up the $CN^{3}\log{N}$ dimension bound supplied by Proposition 3.0.1 for the vector space $R_{2N}$ of all modular functions against the obvious $\gg N^{3}$ lower bound for the subring of the congruence examples from the fact that $\Gamma(2):\Gamma(2N)]\gg N^{3}$ (see Equation 4.2.4). Hence we have:
3.1. A summary
This lays out exactly what we need to prove in order to at least get the unbounded denominators conjecture up to a “small error” (a logarithmic gap $O(\log{N})$ in every level $N$). The following is then a summary of the rest of our paper:
@@ 3.1. A summary
    In §4, we prove that the logarithmic gap between the ring of modular forms with bounded denominators and the ring of congruence modular forms can be leveraged to prove the full unbounded denominators conjecture. The main idea here is that given a noncongruence modular form $f(q)\in\mathbf{Z}\llbracket q^{1/N}\rrbracket$, one can construct many more such forms independent over the congruence ring, by considering $f(q^{1/p})\in\mathbf{Z}\llbracket q^{1/Np}\rrbracket$ for primes $p$.
@@ 4.2. Modular Forms
(The vector space $R_{N}$ was also defined in Proposition 3.0.1 but we repeat the definition here for convenience.) For example, the holomorphic modular forms on $Y(2)$ are given by $\mathbf{Q}(\lambda^{\pm 1},(1-\lambda)^{\pm 1})$ and the $\mathbf{Q}(\lambda)$-vector space generated by such elements inside $\mathbf{Q}(\kern-1.49994pt{(}q^{1/N})\kern-1.49994pt{)}$ is $M_{2}=\mathbf{Q}(\lambda)$.
@@
The second claim follows from Corollary 2.0.5 and the remark that the conformal radius of $\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ is strictly larger than $1$ given in the second paragraph of §1.1.3 (more precisely, $\sqrt[N]{\lambda(z^{N})/16}:D(0,1)\to\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ is a well-defined holomorphic map with derivative being $1$ at 0 and thus the universal covering $D(0,1)\to\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ has derivative strictly larger than $1$ by the Schwarz lemma). ∎
@@
Since $R_{N}$ is finite over $M_{2}$, it is generated by a finite number of basis elements each of which is invariant under some finite index subgroup $G\subset\mathrm{SL}_{2}(\mathbf{Z})$ containing $E$ with $L(G)$ dividing $N$. The intersection of all these groups still has finite index and level $N$ by Lemma 4.1.3, and then we take $G_{N}$ to be the largest normal subgroup of $\mathrm{SL}_{2}(\mathbf{Z})$ contained in this intersection, which also has $L(G_{N})=N$ by Lemma 4.1.4. ∎
Our goal is to prove that $R_{N}=M_{N}$. We can and do assume that $N$ is even. As noted in Lemma 4.2.3, $R_{N}:M_{2}]<\infty$.
The degree of $M_{N}$ over $M_{2}$ is equal to the degree of the modular curve $Y(N)$ over $Y(2)$, and this is given by the explicit formula

### v3

@@ 1. Introduction
We refer the reader to Atkin and Swinnerton-Dyer [ASD71] for the roots of the unbounded denominators conjecture, and to Birch’s article [Bir94] as well as to Long’s survey [Lon08, § 5] for an introduction to this problem and its history. For the vector-valued generalization, see §7.3 and its references below. The cases of relevance to the partition and correlation functions of rational conformal field theories (of which the tip of the iceberg is the example (1.0.3) discussed below) were resolved in a string of works [DR18, DLN15, SZ12, NS10, Xu06, Ban03, Zhu96, AM88], by the modular tensor categories method. Some further sporadic cases of the unbounded denominators conjecture have been settled by mostly ad hoc means [FF22, FM16b, LL12, KL08, KL09].
@@ 1.1.3. A finite local monodromy leads to an overconvergence
But if instead of the tautological simultaneous uniformization we take $\varphi:D(0,1)\to\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ to be the universal covering map (pointed at $\varphi(0)=0$), then either by a direct computation with monodromy, or by Cauchy’s analyticity theorem on the solutions of linear ODEs with analytic coefficients and no singularities in a disc, we have both function germs $x:=\varphi(z)$ and $f(x):=f(\varphi(z))$ holomorphic, hence convergent, on the full unit disc $D(0,1)$. In contrast, now $F(\varphi(z))$ converges only up to the “first” nonzero fiber point in $\varphi^{-1}(0)\smallsetminus\{0\}$, giving a certain radius rather smaller than $1$. We must have the strict lower bound $|\varphi^{\prime}(0)|>1$, because the preceding unit-radius holomorphic map $\sqrt[N]{\lambda(z^{N})/16}:D(0,1)\to\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ has to factorize properly through the universal covering map. Indeed in Theorem 5.1.4, using an explicit description by hypergeometric functions of the multivalued inverse of the universal covering map of $\mathbf{C}\smallsetminus\mu_{N}$ based on Poincaré’s ODE approach [Hem88] to the uniformization of Riemann surfaces, we find an exact formula for this uniformization radius in terms of the Euler Gamma function. [^2] André pointed out to us that this explicit formula has previously been obtained by Kraus and Roth, see [KR16, Remark 5.1]. See also [Gol69, §III.1]. Hence the algebraicity of $f(x)$ gets witnessed by André’s criterion; and the formal new result that we get already at this opening stage (see Theorem 7.2.1) is that any integral formal power series solution $f(x)\in\mathbf{Z}\llbracket x\rrbracket$ to a linear ODE $L(f)=0$ without singularities on $\mathbf{P}^{1}\smallsetminus\{0,1/16,\infty\}$ is in fact algebraic as soon as the linear differential operator $L$ has a finite local monodromy $\mathbf{Z}/N$ around the singular point $x=0$. More than this: the quantitative Corollary 2.0.5 proves that the totality of such $f(x)\in\mathbf{Z}\llbracket x\rrbracket$ at a given $N$ span a finite-dimensional $\mathbf{Q}(x)$-vector space, and gives an upper bound on its dimension as a function of the Wohlfahrt level parameter $N$. Now since a (noncongruence) counterexample $f(\tau)\in\mathbf{Z}\llbracket q\rrbracket$ to Theorem 1.0.1 would not exist on its own but spawn a whole sequence $f(p\tau)\in\mathbf{Z}\llbracket q\rrbracket$ of $\mathbf{Q}(x)$-linearly independent counterexamples at growing Wohlfahrt level $N\mapsto Np$, our idea is to measure up the supply of these putative (fictional) counterexamples alongside the congruence supply at a gradually increasing level until together they break the quantitative bound (2.0.3) supplied by our arithmetic holonomy Theorem 2.0.1.
@@ 1.1.4. The dimension bound can be leveraged with growing level NN
We have the congruence supply of dimension $\Gamma(2):\Gamma(2N)]\gg N^{3}$, and then as a glance at our shape (2.0.7) of holonomy rank bound readily reveals, it seems a fortuitous piece of luck that the conformal size (Riemann uniformization radius at 0) of our relevant Riemann surface $\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ turns out to have the matching asymptotic form $1+\zeta(3)/(2N^{3})+O(N^{-5})$. We “only” have to prove that the numerator (growth) term in the holonomy rank bound (2.0.7) inflates at a slower rate than our extrapolating putative counterexamples $f(\tau)\mapsto f(p\tau)$!
@@ 2. The arithmetic holonomicity theorem
Remark 2.0.10.
We will see in §2.3 that Lemma 2.0.8 is in fact equivalent to Theorem 2.0.1. See however Remark 2.3.3 about a sketch of an improvement.
For a complete proof of the unbounded denominators conjecture, we invite the reader on a first pass to proceed directly to §3 after §2.3.
@@ 2.1. The auxiliary construction
2.2. Extrapolation and proof of Lemma 2.0.8
@@
$\sup_{\mathbf{T}^{d}}\log{|H|}\leq dD\cdot\max\big\{\sup_{\mathbf{T}}\log{|h|},\,\sup_{\mathbf{T}}\log{|h\cdot\varphi^{*}p|}\big\}+\kappa C\alpha+o(\alpha).$
@@
$\alpha\log{|\varphi^{\prime}(0)|}\leq\frac{d}{(d!)^{1/d}}\Big(1+\frac{1}{\kappa}\Big)^{\frac{1}{d}}\cdot\frac{\alpha}{m}\max\big\{\sup_{\mathbf{T}}\log{|h|},\,\sup_{\mathbf{T}}\log{|h\cdot\varphi^{*}p|}\big\}+\kappa C\alpha+o(\alpha)$
@@
$m\leq\inf_{\begin{subarray}{c}d\in\mathbf{N}\\ 0<\kappa<(\log{|\varphi^{\prime}(0)|})/|C|\end{subarray}}\Big\{\frac{\frac{d}{(d!)^{1/d}}\Big(1+\frac{1}{\kappa}\Big)^{\frac{1}{d}}\cdot\max\big\{\sup_{\mathbf{T}}\log{|h|},\,\sup_{\mathbf{T}}\log{|h\cdot\varphi^{*}p|}\big\}}{{\log{|\varphi^{\prime}(0)|-\kappa C}}}\Big\}$
contingent on the denominator being positive. Lemma 2.0.8 now follows by firstly letting $d\to\infty$ and then $\kappa\to 0$, and observing that in that limit
$\frac{d}{(d!)^{1/d}}\Big(1+\frac{1}{\kappa}\Big)^{\frac{1}{d}}\to e\qquad\textrm{ while }\qquad\kappa C\to 0,$
@@
At this point Theorem 2.0.1 comes as the immediate combination of Lemma 2.0.8 and the following classical lemma of Nevanlinna.
@@
Consider a holomorphic function $g:\overline{D(0,1)}\to\mathbf{C}$, and let $\varepsilon>0$. Then there exists a quotient representation
@@
$h(0)=1\quad\textrm{and}\quad\max\big\{\sup_{\mathbf{T}}\log{|h|},\,\sup_{\mathbf{T}}\log{|hg|}\big\}\leq\int_{\mathbf{T}}\log^{+}{|g|}\,\mu_{\mathrm{Haar}}+\varepsilon.$
@@
Remark 2.3.3.
Conversely, for any holomorphic map $h:\overline{D(0,1)}\to\mathbf{C}$ with $h(0)=1$, we have the lower bound
$\max\big\{\sup_{\mathbf{T}}\log{|h|},\,\sup_{\mathbf{T}}\log{|hg|}\big\}\geq\int_{\mathbf{T}}\log^{+}{|g|}\,\mu_{\mathrm{Haar}},$
as one sees immediately from integrating the pointwise identity
$\max\{\log{|h|},\log{|hg|}\}=\log{|h|}+\log^{+}{|g|}$
over $\mathbf{T}$ and using $\int_{\mathbf{T}}\log{|h|}\geq\log{|h(0)|}=0$ from subharmonicity. This shows the necessity of the $\varepsilon$ in Lemma 2.3.1. It also shows that Theorem 2.0.1—our final goal of the current §2, which at this point is fully proved—is in fact equivalent with the intermediate form Lemma 2.0.8.
On the other hand, with a bit more work based on the Law of Large Numbers for restricting all the sums $\sum_{i=1}^{d}k_{i}/D$ involved in Lemma 2.1.2 to be concentrated around the mean $d/2$ under our $d\to\infty$ asymptotic, the proof method of Lemma 2.0.8 does give the finer bound
$m\leq(e/2)\,\inf_{h:\,h(0)=1}\Big\{\frac{\sup_{\mathbf{T}}{\log{|h|}}+\sup_{\mathbf{T}}{\log{|h\cdot\varphi^{*}p|}}}{\log{|\varphi^{\prime}(0)|}}\Big\},$
where the infimum is taken over all holomorphic mappings $h:\overline{D(0,1)}\to\mathbf{C}$ subject to the normalizing constraint $h(0)=1$. We will not need this improvement here.
@@ 2.5.4. A potential-theoretic generalization
In these optics, Bost and Charles [BC22, Corollary 8.3.5] have very recently refined our Theorem 2.0.1 to the cleaner form
$m\leq\frac{\iint_{\mathbf{T}^{2}}\log{|p(\varphi(z))-p(\varphi(w))|}\,\mu_{\mathrm{Haar}}(z)\mu_{\mathrm{Haar}}(w)}{\log{|\varphi^{\prime}(0)|}}.$
In particular, on replacing $p$ by $p^{k}$ with using the elementary inequality $\log{|x-y|}\leq\log^{+}{|x|}+\log^{+}{|y|}+\log{2}$ and taking $k\to+\infty$, their result improves our coefficient $e$ in (2.5.24) to the value $2$. We do not know whether or not this is the best-possible constant.
3. Our approach to the Unbounded Denominators Conjecture
In this section, we lay out our main approach to the unbounded denominators conjecture. This will reduce the proof to a number of independent results in group theory, complex geometry, and complex analysis which we take up in sections §4, §5, and §6. Our main idea is to use our arithmetic holonomicity theorems to prove the following:
@@ 3. Our approach to the Unbounded Denominators Conjecture
We may also prove this proposition by using Theorem 2.0.1 directly. Using the notation in the proof, let $Y^{\prime}$ denote the modular curve $Y$ with all the cusps above $0\in Y(2)\cup\{0\}$ filled in. The fiber product $Y^{\prime}\times_{Y(2)_{\overline{\mathbf{Q}}}\cup\{0\}}U$ with its natural map to $U$ is a covering map (one can check this claim locally; the assumption on cusp widths is used to prove that $0\in U$ is not ramified). Therefore, the universal covering map $D(0,1)\rightarrow U$ factors through $Y^{\prime}\times_{Y(2)_{\overline{\mathbf{Q}}}\cup\{0\}}U$ and thus we obtain a map $D(0,1)\rightarrow Y^{\prime}\times_{Y(2)_{\overline{\mathbf{Q}}}\cup\{0\}}U\rightarrow Y^{\prime}$ (the second map is the natural map) such that its composition with $Y^{\prime}\rightarrow Y(2)_{\overline{\mathbf{Q}}}\cup\{0\}$ is the map $D(0,1)\rightarrow U\rightarrow Y(2)_{\overline{\mathbf{Q}}}\cup\{0\}$. Thus $f\circ\varphi$ is also given by the natural pullback of $f$ from $Y^{\prime}$ to $D(0,1)$ and thus it is holomorphic over $D(0,1)$ as far as $f$ is holomorphic at all cusps in $Y^{\prime}$, which can be achieved by multiplying $f$ with a suitable power of $\lambda$. Thus we verify the analyticity property in Theorem 2.0.1. The rest of the proof is the same as the proof above.
3.1. A guide to the proof of the main theorem
We prove both of the assumptions of Proposition 3.0.1 hold in Theorems 5.1.4 and Theorem 6.0.1 respectively. This provides a $CN^{3}\log{N}$ dimension bound for the vector space $R_{2N}$ of all modular functions against the obvious $\gg N^{3}$ lower bound for the subring of the congruence examples from the fact that $\Gamma(2):\Gamma(2N)]\gg N^{3}$ (see Equation 4.3.3). We then need to provide an additional argument to overcome this “small error” (a logarithmic gap $O(\log{N})$ in every level $N$) between the lower and upper bounds.
The following is a guide to what we do in the next few sections of our paper:
@@ 3.1. A guide to the proof of the main theorem
    In §4, we prove that the logarithmic gap between the ring of modular forms with bounded denominators and the ring of congruence modular forms can be leveraged to prove the full unbounded denominators conjecture. The main idea here is that given a noncongruence modular form $f(q)\in\mathbf{Z}\llbracket q^{1/N}\rrbracket$, one can construct many more such forms independent over the ring of congruence forms by considering $f(q^{p})\in\mathbf{Z}\llbracket q^{1/N}\rrbracket$ for primes $p$.
@@
The following leitfaden gives an abbreviated summary of how the argument is laid out:
@@ 4.2. Modular Forms
(The vector space $R_{N}$ was also defined in Proposition 3.0.1 but we repeat the definition here for convenience.) For example, the (weight 0) holomorphic modular forms on $Y(2)$ are given by $\mathbf{Q}[\lambda,1/\lambda,1/(1-\lambda)]$, and the $\mathbf{Q}(\lambda)$-vector space generated by such elements inside $\mathbf{Q}(\kern-1.49994pt{(}q^{1/N})\kern-1.49994pt{)}$ is $M_{2}=\mathbf{Q}(\lambda)$.
@@
The second claim follows from Corollary 2.0.5 and the remark (cf. the second paragraph of §1.1.3) that the conformal radius of $\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ is strictly larger than $1$. Indeed, $\sqrt[N]{\lambda(z^{N})/16}:D(0,1)\to\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ is a well-defined holomorphic map with unit derivative at the origin, and hence by Schwarz’s lemma the universal covering $D(0,1)\to\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ has derivative strictly larger than $1$ in absolute value. (Later, in Theorem 5.1.4 below, we will exactly compute this latter derivative.) ∎
@@
Since $R_{N}$ is finite over $M_{2}$, it is generated by a finite number of basis elements each of which is invariant under some finite index subgroup $\Phi\subset\mathrm{SL}_{2}(\mathbf{Z})$ containing $E$ with $L(\Phi)$ dividing $N$. The intersection of all these groups still has finite index and level $N$ by Lemma 4.1.3, and then we take $G_{N}$ to be the largest normal subgroup of $\mathrm{SL}_{2}(\mathbf{Z})$ contained in this intersection, which also has $L(G_{N})=N$ by Lemma 4.1.4. ∎
4.3. A leveraging argument
Let us assume that there exists an (even) $N$ such that $R_{N}$ is strictly larger than $M_{N}$. Let $f(\tau)\in\mathbf{Z}\llbracket q^{1/N}\rrbracket\in R_{N}$ be an element which does not lie in $M_{N}$. Recall that all forms in $R_{N}$ and thus in particular $f$ is invariant by a subgroup $G=G_{N}\subset\langle E,\Gamma(N)\rangle$ which is normal with finite index in $\langle E,\Gamma(N)\rangle$ and has $L(G_{N})=N$ by Lemma 4.2.3. The main idea of this section is to exploit the fact that $f(p\tau)\in\mathbf{Z}\llbracket q^{1/N}\rrbracket$ is also a modular form with integer coefficients for any prime $p$. Since the form $f(\tau)$ is invariant under $G$, the form $f(p\tau)$ is invariant under $A^{-1}GA$ and thus also the group $A^{-1}GA\cap\mathrm{SL}_{2}(\mathbf{Z})$. Now, by Lemma 4.1.7, we know that this group has (Wohlfahrt) level dividing $Np$. In particular $f(p\tau)$ has cusp width dividing $Np$ at each cusp, and hence $f(p\tau)\in R_{Np}$.
Our main result is as follows:
Theorem 4.3.1.
Suppose that $(p,N)=1$ is prime. Suppose that $f(\tau)$ is not invariant under a congruence subgroup. Then the form $f(p\tau)$ is not in the $M_{Np}$-algebra generated by $R_{N}$.
That is, we can leverage one exception to the unbounded denominators to produce many examples. Before proving Theorem 4.3.1 (whose proof is deferred to §4.4), we first draw the following consequence:
Theorem 4.3.2.
Suppose that there exists an $N$ such that $R_{N}:M_{N}]>1$. Then for any prime $p$ not dividing $N$, one has
$R_{Np}:M_{Np}]\geq 2[R_{N}:M_{N}].$
Proof.
Let $f(\tau)$ be a form in $R_{Np}$ which is not in $M_{N}$. By Theorem 4.3.1, we deduce that $f(p\tau)\in R_{Np}$ is not in the $M_{Np}$-algebra $R_{N}M_{Np}$ generated by $R_{N}$ (which is a subfield of $R_{Np}$). We have
because the intersection of $M_{Np}$ and $R_{N}$ is $M_{N}$. Thus
$\frac{[R_{Np}:M_{Np}]}{[R_{N}:M_{N}]}=[R_{Np}:R_{N}M_{Np}]$
is an integer which is $\geq 2$, which implies Theorem 4.3.2. ∎
Our goal is to prove that $R_{N}=M_{N}$. We can and do assume that $N$ is even. As noted in Lemma 4.2.3, $R_{N}:M_{2}]<\infty$. The degree of $M_{N}$ over $M_{2}$ is equal to the degree of the modular curve $Y(N)$ over $Y(2)$, and this is given by the explicit formula